OpenStudy (anonymous):

x^2-8x=0 solve by factoring

OpenStudy (saifoo.khan):

x(x-8) = 0 x =0 x -8 = 0 x = 8

OpenStudy (anonymous):

\[x^2-8x=0 \Rightarrow x(x-8) = 0 \Rightarrow x = 0,8 \]

OpenStudy (anonymous):

yw thats correct